Circular No. 6401

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996Z IN NGC 2935 W. Johnson, Anza, CA, reports his discovery of a supernova (mag ∼

16) on a CCD image taken on May 16.17 UT, SN 1996Z located ∼ 1′ south of the center of NGC 2935 (α = 9^{h}36^{m}.7, δ = −21^{o}08′, equinox 2000).

S. Benetti and M. Turatto, European Southern Observatory; and S. Moehler, Space Telescope Science Institute, report: “Inspection of a reduced CCD spectrogram (range 332–924 nm, resolution 2.7 nm) obtained on May 17.97 UT with the ESO/MPI 2.2-m telescope (+ EFOSC2) confirms that this object is a type-Ia supernova. The supernova, as measured on an R CCD frame, is located 4′′ east and 71′′ south of the galaxy nucleus. The spectrum is dominated by P-Cyg lines of intermediate-mass elements (mostly Ca II H and K and infrared triplet; Si II; and S II) superimposed on a relatively blue continuum, which is typical for a type-Ia supernova ∼ 5 days after maximum light. The expansion velocity deduced from the minimum of Si II (635.5 nm), corrected by the redshift (2500 km/s, as measured from a strong Na I D interstellar absorption) of the parent galaxy, is ∼ 12 000 km/s. The mean expansion velocity derived from the Ca II lines is ∼ 16 200 km/s; this is substantially faster than was seen in the prototype-Ia SN 1994D (Patat et al. 1996, MNRAS 278, 111).”

P. Garnavich, A. Riess, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, communicate: “CCD images of NGC 2935 were obtained by R. Jansen on May 18.1 UT using the Whipple Observatory’s 1.2-m telescope. The supernova (V ∼ 14.3) is 70′′ south of the galaxy’s nucleus. A spectrum taken by E. Barton with the 1.5-m Tillinghast telescope on May 18.2 shows that the supernova is of type Ia, a week to ten days past maximum. Strong interstellar Na I (589.3 nm) absorption is detected from our galaxy (equivalent width 0.09 nm) and the host galaxy (equivalent width 0.14 nm).”

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude and coma-diameter estimates (cf. IAUC 6380):

Apr. 14.67 UT, 7.4, ∼ 9′ (D. Seargent, The Entrance, N.S.W., 10×50 binoculars); 18.05, 7.7, 6′ (A. Baransky, Kiev, Ukraine, 0.11-m reflector); 21.48, 7.5, – (C. S. Morris, Pine Mountain Club, CA, 10×50 binoculars); 29.43, 7.8, 7′ (R. Keen, Mt. Thorodin, CO, 0.15-m reflector); May 15.58, 6.7, – (D. Seargent, The Entrance, N.S.W., 2.5×25 binoculars); 16.42, 7.4, 11′ (A. Hale, Cloudcroft, NM, 10×50 binoculars).

1996 May 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green